### v1

# Total variation bounds in the Lindeberg central limit theorem

Abstract
In this paper, we obtain an explicit total variation bound in the central limit theorem for the sums of non-i.i.d. random variables. Our results show that, under suitable assumptions, Lindeberg’s condition is sufficient and necessary for the convergence in total variation distance.
Keywords: Central limit theorem, Total variation distance.
2010 Mathematics Subject Classification: 60F15.
1 Introduction
Let $(X_{k})_{k\geq 1}$ be a sequence of independent real-valued random variables (not necessarily identically distributed) with mean $\mathbb{E}X_{k}=0$ and finite variances $\sigma_{k}^{2}=\mathbb{E}X_{k}^{2}\in(0,\infty).$ We consider the normalized sums
$S_{n}=\frac{X_{1}+\cdots+X_{n}}{b_{n}},$ (1.1)
where $b_{n}=\sqrt{\sigma_{1}^{2}+\sigma_{2}^{2}+\cdots+\sigma_{n}^{2}},\,\,n\geq 1.$
Let us recall the classical central limit theorem: if the random variables $X_{k}$ satisfies Lindeberg’s condition:
$\lim\limits_{n\to\infty}\frac{1}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[X_{k}^{2}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>\varepsilon b_{n}\}}]=0$ (1.2)
for all $\varepsilon>0,$ where $\rm 1\hskip-3.87495pt1$} is the indicator function, then $S_{n}$ converges in distribution to a standard normal random variable NN as $n\to\infty.$ Furthermore, if the random variables $X_{k}$ satisfy the Feller-Lévy condition
$\max\limits_{1\leq k\leq n}(\sigma_{k}^{2}/b_{n}^{2})\to 0\,\,\,\text{as}\,\,n\to\infty,$ (1.3)
then Lindeberg’s condition is both sufficient and necessary. We also have the following optimal bound on the Kolmogorov distance for the rate of convergence (see [3] for a review)
$d_{K}(S_{n},N):=\sup_{x\in\mathbb{R}}|P(S_{n}\leq x)-P(N\leq x)|$ $\leq\frac{c}{b_{n}^{3}}\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)]$ (1.4)
$\leq\frac{c}{b_{n}^{3}}\sum\limits_{k=1}^{n}\mathbb{E}|X_{k}|^{3},$ (1.5)
where $c$ is a positive absolute constant. We now consider the total variation distance between the laws of $S_{n}$ and NN defined by
$d_{TV}(S_{n},N):=\frac{1}{2}\sup_{\|h\|_{\infty}\leq 1}|Eh(S_{n})-Eh(N)|,$
where $h$ is measurable and $\|h\|_{\infty}:=\sup\limits_{x\in\mathbb{R}}|h(x)|.$ Note that the convergence in total variation is stronger than the convergence in distribution. The study of the central limit theorem in total variation distance has a long history beginning in 1952 with the results of Prokhorov [12]. However, the problem of obtaining optimal estimates for $d_{TV}(S_{n},N)$ (under minimal moment conditions) is non-trivial. The reader can consults [1, 13] and references therein for the case of independent identically distributed (i.i.d.) random variables with finite third absolute moment. For the case of non-i.i.d. random variables, it is natural to wonder whether the bounds (1.4) and (1.5) may be strengthened to the total variation distance as well.
When all summands $S_{n}$ have discrete distributions, $d_{TV}(S_{n},N)=1\nrightarrow 0$ as $n\to\infty.$ Hence, to obtain an affirmative answer, we need to impose additional assumptions on the sequence $(X_{k})_{k\geq 1}.$ We have the following result, due to Bobkov-Chistyakov-Götze [5]: Assume that the independent random variables $X_{1},\cdots,X_{n}$ have finite third absolute moments, and that $D(X_{k})\leq D$ for all $1\leq k\leq n$ and for some $D>0.$ Then
$d_{TV}(S_{n},N)\leq\frac{C_{D}}{b_{n}^{3}}\sum\limits_{k=1}^{n}\mathbb{E}|X_{k}|^{3},$
where $C_{D}$ is a positive constant depending only on $D.$ Here $D(.)$ denotes the relative entropy defined as follows: for a random variable $X$ with density $p(x)$ and finite variance,
$D(X)=\int_{-\infty}^{\infty}p(x)\log\frac{p(x)}{\varphi_{X}(x)}dx,$ where $\varphi_{X}(x)=\frac{1}{\sqrt{2\pi{\rm Var}(X)}}e^{-\frac{(x-\mathbb{E}X)^{2}}{2{\rm Var}(X)}},x\in\mathbb{R}.$
The aim of this paper is to establish the bound (1.4) for $d_{TV}(S_{n},N).$ Standing for the finiteness of relative entropies, we require the random variables $X_{k}^{\prime}s$ to have the finite standardised Fisher information. We recall that, given a random variable $X$ with an absolutely continuous density $p(x),$ the standardised Fisher information of $X$ (or its distribution) is defined by
$J_{st}(X)={\rm Var}(X)\int_{-\infty}^{+\infty}\frac{p^{\prime}(x)^{2}}{p(x)}dx-1={\rm Var}(X)\mathbb{E}[\rho^{2}(X)]-1,$
where $p^{\prime}$ denotes a Radon-Nikodym derivative of $p$ and $\rho:=p^{\prime}/p$ is the score function. For more details about the Fisher information and the central limit theorem results, we refer the reader to the monograph [9] and recent papers [6, 8, 10].
Hereafter, we put $J(X)=J_{st}(X)+1.$ Our main results are stated in the following theorem.
Theorem $1.$
Let $(X_{k})_{k\geq 1}$ be a sequence of independent real-valued random variables $\mathbb{E}X_{k}=0$ and finite variances $\sigma_{k}^{2}=\mathbb{E}X_{k}^{2}\in(0,\infty).$ Assume that, for every $k\geq 1,$ the law of $X_{k}$ has an absolutely continuous density $p_{k}$ satisfying $J(X_{k})<\infty.$ Then, for $S_{n}$ defined by (1.1), we have
$d_{TV}(S_{n},N)\leq\bigg(\frac{8\pi\max\limits_{1\leq k\leq n}J(X_{k})}{1-\max\limits_{1\leq k\leq n}(\sigma_{k}^{2}/b_{n}^{2})}\bigg)^{\frac{1}{2}}\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)]}{b_{n}^{3}}.$ (1.6)
Our condition $J(X_{k})<\infty$ is stronger than the condition $D(X_{k})<\infty$ required in [5]. Indeed, we have $D(X_{k})\leq(J(X_{k})-1)/2,$ see e.g. Lemma E.3 in [9]. However, our bound (1.6) allows us to upgrade the Lindeberg central limit theorem to the convergence in total variation distance as follows.
Corollary $1.$
Let the sequence $(X_{k})_{k\geq 1}$ be as in Theorem $1.$ Assume additionally that $\sup\limits_{k\geq 1}J(X_{k})<\infty.$ The, the following two statements are equivalent
$(i)$ Lindeberg’s condition (1.2) is satisfied.
$(ii)$ $d_{TV}(S_{n},N)\to 0$ as $n\to\infty$ and the Feller-Lévy condition (1.3) is satisfied.
2 Proofs
We first prepare some technical results. For a given real valued measurable function $h$ with $\mathbb{E}h(Z)<\infty.$ The ordinary differential equation
$f^{\prime}(x)-xf(x)=h(x)-\mathbb{E}h(Z)$ (2.1)
is called Stein’s equation associated with $h$. The unique bounded solution of (2.1) is given by
$f(x):=e^{x^{2}/2}\int_{-\infty}^{x}(h(y)-\mathbb{E}h(Z))e^{-y^{2}/2}dy,\,\,x\in\mathbb{R}.$ (2.2)
We recall from Lemma 2.4 in [7] the following estimates for $f.$
Lemma 2.1.
Let $f$ be the solution (2.2) to Stein’s equation. If $h$ is bounded $1,$ then
$\text{$\|f\|_{\infty}\leq\sqrt{2\pi}$ and $\|f^{\prime}\|_{\infty}\leq 4$}.$
Lemma $2.$
Let the sequence $(X_{k})_{k\geq 1}$ be as in Theorem $1.$ For each $1\leq k\leq n,$ we put
$S_{k,n}=\frac{X_{1}+\cdots+X_{k-1}+X_{k+1}+\cdots+X_{n}}{b_{k,n}},$
where $b_{k,n}=\sqrt{b_{n}^{2}-\sigma_{k}^{2}}.$ Then the law of $S_{k,n}$ has an absolutely continuous density. Moreover, the score function $\rho_{k,n}$ of $S_{k,n}$ is given by
$\rho_{k,n}(x)=\mathbb{E}\left[\frac{\sum\limits_{1\leq i\leq n,i\neq k}\sigma_{i}^{2}\rho_{i}(X_{i})}{b_{k,n}}\bigg|S_{k,n}=x\right]$ (2.3)
Proof.
For each $k\geq 1,$ it follows from Theorem VIII.1 in [2] that the density $p_{k}$ of the random variable $X_{k}$ satisfies
$\lim\limits_{x\to\pm\infty}(1+|x|)p_{k}(x)=0.$
Hence, for any bounded and differentiable function $f:\mathbb{R}\to\mathbb{R}$ with bounded derivative, we can use Lemma 1.20 in [9] to get
$\mathbb{E}[f(X_{k})\rho_{k}(X_{k})]=-\mathbb{E}[f^{\prime}(X_{k})].$
Consequently, by the independence of $X_{k}^{\prime}s,$ we obtain
$\mathbb{E}\left[f(S_{k,n})\frac{\sum\limits_{1\leq i\leq n,i\neq k}\sigma_{i}^{2}\rho_{i}(X_{i})}{b_{k,n}}\right]$ $=\frac{\sum\limits_{1\leq i\leq n,i\neq k}\sigma_{i}^{2}\mathbb{E}\left[f(S_{k,n})\rho_{i}(X_{i})\right]}{b_{k,n}}$
$=-\mathbb{E}[f^{\prime}(S_{k,n})].$ (2.4)
Then, by using the same arguments as in the proof of Proposition 3.2 in [8], we conclude that the law of $S_{k,n}$ has an absolutely continuous density and its score function is given by (2.3). This finishes the proof of the lemma. ∎
Definition 2.1.
Given a random variable $Y\in L^{1}(\Omega)$ with a density $p_{Y}$ supported on some connected interval of $\mathbb{R},$ we define the Stein kernel $\tau_{Y}$ associated to $Y$ as follows
$\tau_{Y}(x)=\frac{\int_{x}^{\infty}(y-\mathbb{E}[Y])p_{Y}(y)dy}{p_{Y}(x)}$
for $x$ in the support of $p_{Y}.$
In the proofs, we use the following fundamental properties of Stein kernels, see e.g. [11] and references therein: $\tau_{Y}(Y)\geq 0,$ $\mathbb{E}[\tau_{Y}(Y)]={\rm Var}(Y)$ and
$\mathbb{E}[f(Y)(Y-\mathbb{E}Y)]=\mathbb{E}[f^{\prime}(Y)\tau_{Y}(Y)]$ (2.5)
for any differentiable function $f:\mathbb{R}\to\mathbb{R}$ with bounded derivative.
We now are ready to prove the main results of the present paper.
Proof of Theorem $1.$ We will carry out the proof in two steps.
Step 1. In this step, we additionally assume that
For every $k\geq 1,$ the density $p_{k}$ of $X_{k}$ is supported on some connected interval of $\mathbb{R}.$ (2.6)
According to Definition 2.1, this assumption allows us to define the Stein kernels
$\tau_{k}(X_{k})=\frac{\int_{X_{k}}^{\infty}yp_{k}(y)dy}{p_{k}(x)},\,\,k\geq 1.$
Let $h$ be a measurable function bounded by $1.$ From the Stein equation (2.1) we have
$\mathbb{E}h(S_{n})-\mathbb{E}h(N)$ $=\mathbb{E}f^{\prime}(S_{n})-\mathbb{E}[S_{n}f(S_{n})]$
$=\mathbb{E}f^{\prime}(S_{n})-\frac{1}{b_{n}}\sum\limits_{k=1}^{n}\mathbb{E}[X_{k}f(S_{n})].$
Hence, by the property (2.5) of the Stein kernels, we deduce
$\mathbb{E}h(S_{n})-\mathbb{E}h(N)$ $=\mathbb{E}f^{\prime}(S_{n})-\frac{1}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[\tau_{k}(X_{k})f^{\prime}(S_{n})]$
$=\frac{1}{b_{n}^{2}}\mathbb{E}\left[f^{\prime}(S_{n})(b_{n}^{2}-\tau_{1}(X_{1})-\cdots-\tau_{n}(X_{n}))\right]$
$=\frac{1}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}\left[f^{\prime}(S_{n})(\sigma_{k}^{2}-\tau_{k}(X_{k}))\right].$ (2.7)
Fixed $1\leq k\leq n.$ Let $S_{k,n}$ be as in Lemma $2.$ By the independence of $X_{k}$ and $S_{k,n},$ we can use the relation (2.4) to get
$\mathbb{E}\left[f^{\prime}(S_{n})(\sigma_{k}^{2}-\tau_{k}(X_{k}))\right]$ $=\mathbb{E}\left[f^{\prime}\bigg(\frac{b_{k,n}S_{k,n}+X_{k}}{b_{n}}\bigg)(\sigma_{k}^{2}-\tau_{k}(X_{k}))\right]$
$=-\frac{b_{n}}{b_{k,n}}\mathbb{E}\left[f\bigg(\frac{b_{k,n}S_{k,n}+X_{k}}{b_{n}}\bigg)(\sigma_{k}^{2}-\tau_{k}(X_{k}))\rho_{k,n}(S_{k,n})\right].$
On the other hand, since $\mathbb{E}[\tau_{k}(X_{k})]=\sigma_{k}^{2},$ we have
$\mathbb{E}\left[f\bigg(\frac{b_{k,n}S_{k,n}}{b_{n}}\bigg)(\sigma_{k}^{2}-\tau_{k}(X_{k}))\rho_{k,n}(S_{k,n})\right]=\mathbb{E}\left[f\bigg(\frac{b_{k,n}S_{k,n}}{b_{n}}\bigg)\rho_{k,n}(S_{k,n})\right]\mathbb{E}[\sigma_{k}^{2}-\tau_{k}(X_{k})]=0.$
So we obtain
$\mathbb{E}\left[f^{\prime}(S_{n})(\sigma_{k}^{2}-\tau_{k}(X_{k}))\right]$
$=-\frac{b_{n}}{b_{k,n}}\mathbb{E}\left[\bigg(f\bigg(\frac{b_{k,n}S_{k,n}+X_{k}}{b_{n}}\bigg)-f\bigg(\frac{b_{k,n}S_{k,n}}{b_{n}}\bigg)\bigg)(\sigma_{k}^{2}-\tau_{k}(X_{k}))\rho_{k,n}(S_{k,n})\right].$ (2.8)
Furthermore, we have
$\bigg|f\bigg(\frac{b_{k,n}S_{k,n}+X_{k}}{b_{n}}\bigg)-f\bigg(\frac{b_{k,n}S_{k,n}}{b_{n}}\bigg)\bigg|$
$=\bigg|f\bigg(\frac{b_{k,n}S_{k,n}+X_{k}}{b_{n}}\bigg)-f\bigg(\frac{b_{k,n}S_{k,n}}{b_{n}}\bigg)\bigg|{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>b_{n}\}}+\bigg|f\bigg(\frac{b_{k,n}S_{k,n}+X_{k}}{b_{n}}\bigg)-f\bigg(\frac{b_{k,n}S_{k,n}}{b_{n}}\bigg)\bigg|{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|\leq b_{n}\}}.$
Then, by using the estimates $\|f\|_{\infty}\leq\sqrt{2\pi}$ and $\|f^{\prime}\|_{\infty}\leq 4$ from Lemma 2.1, we deduce
$\bigg|f\bigg(\frac{b_{k,n}S_{k,n}+X_{k}}{b_{n}}\bigg)-f\bigg(\frac{b_{k,n}S_{k,n}}{b_{n}}\bigg)\bigg|$ $\leq\sqrt{8\pi}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>b_{n}\}}+\frac{4|X_{k}|}{b_{n}}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|\leq b_{n}\}}$
$\leq\frac{\sqrt{8\pi}}{b_{n}}(b_{n}\wedge|X_{k}|).$ (2.9)
Inserting (2.9) into (2.8) yields
$|\mathbb{E}$ $\left[f^{\prime}(S_{n})(\sigma_{k}^{2}-\tau_{k}(X_{k}))\right]|\leq\frac{\sqrt{8\pi}}{b_{k,n}}\mathbb{E}\left[(b_{n}\wedge|X_{k}|)|(\sigma_{k}^{2}-\tau_{k}(X_{k}))\rho_{k,n}(S_{k,n})|\right]$
$\leq\frac{\sqrt{8\pi}}{b_{k,n}}\mathbb{E}\left[(b_{n}\wedge|X_{k}|)|\sigma_{k}^{2}-\tau_{k}(X_{k})|\right]\mathbb{E}|\rho_{k,n}(S_{k,n})|$
$\leq\frac{\sqrt{8\pi}}{b_{k,n}}\mathbb{E}|\rho_{k,n}(S_{k,n})|\mathbb{E}\left[(b_{n}\wedge|X_{k}|)\tau_{k}(X_{k})\right]+\frac{\sqrt{8\pi}}{b_{k,n}}\mathbb{E}|\rho_{k,n}(S_{k,n})|\mathbb{E}\left[(b_{n}\wedge|X_{k}|)\sigma_{k}^{2}\right].$ (2.10)
Once again, we use the property (2.5) of the Stein kernels to get the following estimate for the first addend in the right hand side of (2.10)
$\mathbb{E}[(b_{n}\wedge|X_{k}|)\tau_{k}(X_{k})]$ $=\mathbb{E}\left[X_{k}\int_{0}^{X_{k}}(b_{n}\wedge|y|)dy\right]$
$\leq\mathbb{E}\left[|X_{k}|\int_{0}^{|X_{k}|}(b_{n}\wedge y)dy\right]$
$\leq\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)].$
For the second addend in the right hand side of (2.10), by Chebyshev’s association inequality, we have
$\mathbb{E}[(b_{n}\wedge|X_{k}|)\sigma_{k}^{2}]=\mathbb{E}[(b_{n}\wedge|X_{k}|)]\mathbb{E}[|X_{k}|^{2}]\leq\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)].$
So it holds that
$|\mathbb{E}\left[f^{\prime}(S_{n})(\sigma_{k}^{2}-\tau_{k}(X_{k}))\right]|\leq\frac{2\sqrt{8\pi}}{b_{k,n}}\mathbb{E}|\rho_{k,n}(S_{k,n})|\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)],\,\,1\leq k\leq n.$
From the above estimates and (2.7), we conclude that
$|\mathbb{E}h(S_{n})-\mathbb{E}h(N)|\leq\frac{2\sqrt{8\pi}}{b_{n}^{2}}\sum\limits_{k=1}^{n}\frac{\mathbb{E}|\rho_{k,n}(S_{k,n})|\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)]}{b_{k,n}}$
for any measurable function $h$ bounded by $1.$
Step 2. In this step, we remove the assumption (2.6) and conclude the proof. Let $(N_{k})_{k\geq 1}$ be a sequence of independent standard normal random variables, and independence of $(X_{k})_{k\geq 1}.$ Fixed $\delta>0,$ we put $\tilde{X}_{k}=X_{k}+\delta N_{k}$ and
$\tilde{S}_{n}=\frac{\tilde{X}_{1}+\cdots+\tilde{X}_{n}}{\tilde{b}_{n}},\,\,\,\tilde{S}_{k,n}=\frac{\tilde{X}_{1}+\cdots+\tilde{X}_{k-1}+\tilde{X}_{k+1}+\cdots+\tilde{X}_{n}}{\tilde{b}_{k,n}},$
where $\tilde{b}_{n}=\sqrt{\sigma_{1}^{2}+\sigma_{2}^{2}+\cdots+\sigma_{n}^{2}+n\delta^{2}}$ and $\tilde{b}_{k,n}=\sqrt{\tilde{b}_{n}^{2}-\sigma_{k}^{2}-\delta^{2}}.$ Since the sequence $(\tilde{X}_{k})_{k\geq 1}$ fulfils the assumption (2.6), this allows to use the result proved in Step 1 and we obtain
$|\mathbb{E}h(\tilde{S}_{n})-\mathbb{E}h(N)|\leq\frac{2\sqrt{8\pi}}{\tilde{b}_{n}^{2}}\sum\limits_{k=1}^{n}\frac{\mathbb{E}|\tilde{\rho}_{k,n}(\tilde{S}_{k,n})|\mathbb{E}[|\tilde{X}_{k}|^{2}(\tilde{b}_{n}\wedge|\tilde{X}_{k}|)]}{\tilde{b}_{k,n}}$ (2.11)
for any measurable function $h$ bounded by $1,$ where $\tilde{\rho}_{k,n}$ denotes the score function of $\tilde{S}_{k,n}.$
Obviously, we have $\lim\limits_{\delta\to 0}\tilde{b}_{n}^{2}=b_{n}^{2},$ $\lim\limits_{\delta\to 0}\tilde{b}_{k,n}=b_{k,n}$ and $\lim\limits_{\delta\to 0}\mathbb{E}[|\tilde{X}_{k}|^{2}(\tilde{b}_{n}\wedge|\tilde{X}_{k}|)]=\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)].$ We also have $\lim\limits_{\delta\to 0}\mathbb{E}|\tilde{\rho}_{k,n}(\tilde{S}_{k,n})|)=\mathbb{E}|\rho_{k,n}(S_{k,n})|$ by Theorem 1.4 in [4]. As a consequence, by letting $\delta\to 0,$ we obtain from (2.11) that
$|\mathbb{E}h(S_{n})-\mathbb{E}h(N)|\leq\frac{2\sqrt{8\pi}}{b_{n}^{2}}\sum\limits_{k=1}^{n}\frac{\mathbb{E}|\rho_{k,n}(S_{k,n})|\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)]}{b_{k,n}}$
for any continuous function $h$ bounded by $1.$ In addition, from the representation formula (2.3), we have
$\mathbb{E}|\rho_{k,n}(S_{k,n})|$ $\leq\sqrt{\mathbb{E}|\rho_{k,n}(S_{k,n})|^{2}}\leq\frac{1}{b_{k,n}}\bigg(\mathbb{E}\bigg|\sum\limits_{1\leq i\leq n,i\neq k}\sigma_{i}^{2}\rho_{i}(X_{i})\bigg|^{2}\bigg)^{\frac{1}{2}}$
$=\frac{1}{b_{k,n}}\bigg(\sum\limits_{1\leq i\leq n,i\neq k}\sigma_{i}^{4}\mathbb{E}|\rho_{i}(X_{i})|^{2}\bigg)^{\frac{1}{2}}=\frac{1}{b_{k,n}}\bigg(\sum\limits_{1\leq i\leq n,i\neq k}\sigma_{i}^{2}J(X_{i})\bigg)^{\frac{1}{2}}$
$\leq\big(\max\limits_{1\leq k\leq n}J(X_{k})\big)^{\frac{1}{2}}.$
We therefore obtain, for any continuous function $h$ bounded by $1,$
$|\mathbb{E}h(S_{n})-\mathbb{E}h(N)|$ $\leq\frac{2\big(8\pi\max\limits_{1\leq k\leq n}J(X_{k})\big)^{\frac{1}{2}}}{b_{n}^{2}}\sum\limits_{k=1}^{n}\frac{\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)]}{b_{k,n}}$
$=\frac{2\big(8\pi\max\limits_{1\leq k\leq n}J(X_{k})\big)^{\frac{1}{2}}}{b_{n}^{3}}\sum\limits_{k=1}^{n}\frac{\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)]}{\big(1-\sigma_{k}^{2}/b_{n}\big)^{\frac{1}{2}}}$
$\leq 2\bigg(\frac{8\pi\max\limits_{1\leq k\leq n}J(X_{k})}{1-\max\limits_{1\leq k\leq n}(\sigma_{k}^{2}/b_{n}^{2})}\bigg)^{\frac{1}{2}}\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)]}{b_{n}^{3}}.$ (2.12)
By an application of Lusin’s theorem,
$d_{TV}(S_{n},N)=\frac{1}{2}\sup_{h\in\mathcal{C},\|h\|_{\infty}\leq 1}|Eh(S_{n})-Eh(N)|,$
where $\mathcal{C}$ denotes the space of continuous functions. So the bound (1.6) follows directly from (2.12). The proof of Theorem 1.1 is complete. $\square$
Proof of Corollary $1.$ The implication $(ii)\Rightarrow(i)$ is obvious because of the fact that the convergence in total variation implies the convergence in distribution. To verify the implication $(i)\Rightarrow(ii),$ we observe that, for any $\varepsilon\in(0,1),$ we have
$\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)]}{b_{n}^{3}}=\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>b_{n}\}}]}{b_{n}^{2}}+\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{3}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|\leq b_{n}\}}]}{b_{n}^{3}}$
$=\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>b_{n}\}}]}{b_{n}^{2}}+\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{3}{\rm 1\hskip-3.87495pt1}_{\{\varepsilon b_{n}<|X_{k}|\leq b_{n}\}}]}{b_{n}^{3}}+\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{3}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|\leq\varepsilon b_{n}\}}]}{b_{n}^{3}}$
$\leq\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>b_{n}\}}]}{b_{n}^{2}}+\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{\varepsilon b_{n}<|X_{k}|\leq b_{n}\}}]}{b_{n}^{2}}+\frac{\varepsilon\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{3}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|\leq\varepsilon b_{n}\}}]}{b_{n}^{2}}$
$\leq\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>\varepsilon b_{n}\}}]}{b_{n}^{2}}+\varepsilon.$
On the other hand, it is well known that Lindeberg’s condition implies the Feller-Lévy condition (1.3). Hence, recalling the bound (1.6), we deduce
$\lim\limits_{n\to\infty}d_{TV}(S_{n},N)\leq\big(8\pi\sup\limits_{k\geq 1}J(X_{k})\big)^{\frac{1}{2}}\varepsilon.$
Letting $\varepsilon\to 0$ gives us the desired conclusion. $\square$

### v2

@@
 $\lim\limits_{n\to\infty}\frac{1}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[X_{k}^{2}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>\varepsilon b_{n}\}}]=0$ (1.2)
-for all $\varepsilon>0,$ where $\rm 1\hskip-3.87495pt1$} is the indicator function, then $S_{n}$ converges in distribution to a standard normal random variable NN as $n\to\infty.$ Furthermore, if the random variables $X_{k}$ satisfy the Feller-Lévy condition
+for all $\varepsilon>0,$ where $\rm 1\hskip-3.87495pt1$} is the indicator function, then $S_{n}$ converges in distribution to a standard normal random variable $N$ as $n\to\infty.$ Furthermore, if the random variables $X_{k}$ satisfy the Feller-Lévy condition
 $\max\limits_{1\leq k\leq n}(\sigma_{k}^{2}/b_{n}^{2})\to 0\,\,\,\text{as}\,\,n\to\infty,$ (1.3)
@@
 $\leq\frac{c}{b_{n}^{3}}\sum\limits_{k=1}^{n}\mathbb{E}|X_{k}|^{3},$ (1.5)
-where $c$ is a positive absolute constant. We now consider the total variation distance between the laws of $S_{n}$ and NN defined by
+where $c$ is a positive absolute constant. We now consider the total variation distance between the laws of $S_{n}$ and $N$ defined by
 $d_{TV}(S_{n},N):=\frac{1}{2}\sup_{\|h\|_{\infty}\leq 1}|Eh(S_{n})-Eh(N)|,$
@@
 Hereafter, we put $J(X)=J_{st}(X)+1.$ Our main results are stated in the following theorem.
-Theorem $1.$
+Theorem 1.1.
 Let $(X_{k})_{k\geq 1}$ be a sequence of independent real-valued random variables $\mathbb{E}X_{k}=0$ and finite variances $\sigma_{k}^{2}=\mathbb{E}X_{k}^{2}\in(0,\infty).$ Assume that, for every $k\geq 1,$ the law of $X_{k}$ has an absolutely continuous density $p_{k}$ satisfying $J(X_{k})<\infty.$ Then, for $S_{n}$ defined by (1.1), we have
@@
 However, our bound (1.6) allows us to upgrade the Lindeberg central limit theorem to the convergence in total variation distance as follows.
-Corollary $1.$
-Let the sequence $(X_{k})_{k\geq 1}$ be as in Theorem $1.$ Assume additionally that $\sup\limits_{k\geq 1}J(X_{k})<\infty.$ The, the following two statements are equivalent
+Corollary 1.1.
+Let the sequence $(X_{k})_{k\geq 1}$ be as in Theorem 1.1. Assume additionally that $\sup\limits_{k\geq 1}J(X_{k})<\infty.$ The, the following two statements are equivalent
 $(i)$ Lindeberg’s condition (1.2) is satisfied.
 $(ii)$ $d_{TV}(S_{n},N)\to 0$ as $n\to\infty$ and the Feller-Lévy condition (1.3) is satisfied.
+We use the condition $J(X_{k})<\infty$ is to establish the bound (1.6). However, this condition can be removed to recover the classical central limit theorem.
+Corollary 1.2.
+Let $(Y_{k})_{k\geq 1}$ be a sequence of independent real-valued random variables $\mathbb{E}Y_{k}=0$ and finite variances $\alpha_{k}^{2}=\mathbb{E}Y_{k}^{2}\in(0,\infty).$ Assume that this sequence satisfies Lindeberg’s condition:
+$\lim\limits_{n\to\infty}\frac{1}{a_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[Y_{k}^{2}{\rm 1\hskip-3.87495pt1}_{\{|Y_{k}|>\varepsilon a_{n}\}}]=0$ (1.7)
+for all $\varepsilon>0,$ where $a_{n}:=\sqrt{\alpha_{1}^{2}+\cdots+\alpha_{n}^{2}}.$ Then, as $n\to\infty,$ we have $\frac{Y_{1}+\cdots+Y_{n}}{a_{n}}\to N$ in distribution.
 2 Proofs
@@
 $\text{$\|f\|_{\infty}\leq\sqrt{2\pi}$ and $\|f^{\prime}\|_{\infty}\leq 4$}.$
-Lemma $2.$
-Let the sequence $(X_{k})_{k\geq 1}$ be as in Theorem $1.$ For each $1\leq k\leq n,$ we put
+Lemma 2.2.
+Let the sequence $(X_{k})_{k\geq 1}$ be as in Theorem 1.1. For each $1\leq k\leq n,$ we put
 $S_{k,n}=\frac{X_{1}+\cdots+X_{k-1}+X_{k+1}+\cdots+X_{n}}{b_{k,n}},$
 where $b_{k,n}=\sqrt{b_{n}^{2}-\sigma_{k}^{2}}.$ Then the law of $S_{k,n}$ has an absolutely continuous density. Moreover, the score function $\rho_{k,n}$ of $S_{k,n}$ is given by
-$\rho_{k,n}(x)=\mathbb{E}\left[\frac{\sum\limits_{1\leq i\leq n,i\neq k}\sigma_{i}^{2}\rho_{i}(X_{i})}{b_{k,n}}\bigg|S_{k,n}=x\right]$ (2.3)
+$\rho_{k,n}(x)=\mathbb{E}\left[\frac{\sum\limits_{1\leq i\leq n,i\neq k}\sigma_{i}^{2}\rho_{i}(X_{i})}{b_{k,n}}\bigg|S_{k,n}=x\right],$ (2.3)
+where $\rho_{i}$ denotes the score function of $X_{i}.$
 Proof.
@@
 $=-\mathbb{E}[f^{\prime}(S_{k,n})].$ (2.4)
-Then, by using the same arguments as in the proof of Proposition 3.2 in [8], we conclude that the law of $S_{k,n}$ has an absolutely continuous density and its score function is given by (2.3).
+Then, by using the same arguments as in the proof of Proposition 3.2 in our recent paper [8], we conclude that the law of $S_{k,n}$ has an absolutely continuous density and its score function is given by (2.3).
 This finishes the proof of the lemma.
@@
 We now are ready to prove the main results of the present paper.
-Proof of Theorem $1.$ We will carry out the proof in two steps.
+Proof of Theorem 1.1. We will carry out the proof in two steps.
 Step 1. In this step, we additionally assume that
@@
 $=\frac{1}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}\left[f^{\prime}(S_{n})(\sigma_{k}^{2}-\tau_{k}(X_{k}))\right].$ (2.7)
-Fixed $1\leq k\leq n.$ Let $S_{k,n}$ be as in Lemma $2.$ By the independence of $X_{k}$ and $S_{k,n},$ we can use the relation (2.4) to get
+Fixed $1\leq k\leq n.$ Let $S_{k,n}$ be as in Lemma 2.2. By the independence of $X_{k}$ and $S_{k,n},$ we can use the relation (2.4) to get
 $\mathbb{E}\left[f^{\prime}(S_{n})(\sigma_{k}^{2}-\tau_{k}(X_{k}))\right]$ $=\mathbb{E}\left[f^{\prime}\bigg(\frac{b_{k,n}S_{k,n}+X_{k}}{b_{n}}\bigg)(\sigma_{k}^{2}-\tau_{k}(X_{k}))\right]$
@@
 where $\mathcal{C}$ denotes the space of continuous functions. So the bound (1.6) follows directly from (2.12). The proof of Theorem 1.1 is complete. $\square$
-Proof of Corollary $1.$ The implication $(ii)\Rightarrow(i)$ is obvious because of the fact that the convergence in total variation implies the convergence in distribution.
+Proof of Corollary 1.1.
+The implication $(ii)\Rightarrow(i)$ is obvious because of the fact that the convergence in total variation implies the convergence in distribution.
 To verify the implication $(i)\Rightarrow(ii),$ we observe that, for any $\varepsilon\in(0,1),$ we have
@@
 Letting $\varepsilon\to 0$ gives us the desired conclusion. $\square$
+Proof of Corollary 1.2.
+Let $(N_{k})_{k\geq 1}$ be a sequence of independent normal random variables, and independence of $(Y_{k})_{k\geq 1}.$ For each $k\geq 1,$ $\mathbb{E}N_{k}=0$ and $\mathbb{E}N_{k}^{2}=\mathbb{E}Y_{k}^{2}.$ We consider the sequence $X_{k}=Y_{k}+N_{k},k\geq 1$ and define $S_{n}$ as in (1.1).
+For every $\varepsilon>0,$ we have
+$\frac{1}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>\varepsilon b_{n}\}}]$ $=\frac{1}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[|Y_{k}+N_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|Y_{k}+N_{k}|>\varepsilon b_{n}\}}]$
+$\leq\frac{2}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[(|Y_{k}|^{2}+|N_{k}|^{2}){\rm 1\hskip-3.87495pt1}_{\{|Y_{k}|+|N_{k}|>\varepsilon b_{n}\}}]$
+$\leq\frac{2}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[(|Y_{k}|^{2}+|N_{k}|^{2})({\rm 1\hskip-3.87495pt1}_{\{|Y_{k}|>\varepsilon b_{n}/2\}}+{\rm 1\hskip-3.87495pt1}_{\{|N_{k}|>\varepsilon b_{n}/2\}})].$
+By Chebyshev’s association inequality we deduce
+$\frac{1}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>\varepsilon b_{n}\}}]$ $\leq\frac{4}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[|Y_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|Y_{k}|>\varepsilon b_{n}/2\}}]+\frac{4}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[|N_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|N_{k}|>\varepsilon b_{n}/2\}}].$
+It is easy to check that $\frac{4}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[|N_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|N_{k}|>\varepsilon b_{n}/2\}}]\to 0$ as $n\to\infty.$ Hence, by the condition (1.7), we conclude that the sequence $(X_{k})_{k\geq 1}$ also satisfies Lindeberg’s condition.
+On the other hand, by Lemma 1.20 in [9], the score function of $X_{k}$ is given by
+$\rho_{k}(X_{k})=\mathbb{E}\left[\frac{-N_{k}}{\alpha_{k}^{2}}\bigg|X_{k}\right],\,\,k\geq 1.$
+This implies that $\sup\limits_{k\geq 1}J(X_{k})<\infty.$ Indeed, we have
+$J(X_{k})={\rm Var}(X_{k})\mathbb{E}[\rho_{k}^{2}(X_{k})]=2\alpha_{k}^{2}\mathbb{E}\left(\mathbb{E}\left[\frac{-N_{k}}{\alpha_{k}^{2}}\bigg|X_{k}\right]\right)^{2}\leq 2,\,\,k\geq 1.$
+Thanks to Corollary 1.1, we obtain $d_{TV}(S_{n},N)\to 0$ as $n\to\infty.$ For every $t\in\mathbb{R},$ in view of Lévy’s continuity theorem, we have
+$\mathbb{E}\left[\exp\left(it\frac{Y_{1}+\cdots+Y_{n}}{a_{n}}\right)\right]$ $=e^{t^{2}/2}\mathbb{E}\left[\exp\left(it\frac{Y_{1}+N_{1}+\cdots+Y_{n}+N_{n}}{a_{n}}\right)\right]$
+$=e^{t^{2}/2}\mathbb{E}\left[\exp\left(it\sqrt{2}S_{n}\right)\right]\to e^{-t^{2}/2},\,\,n\to\infty$
+and hence, $\frac{Y_{1}+\cdots+Y_{n}}{a_{n}}\to N$ in distribution.
+The proof is complete. $\square$